One mole of an ideal gas is heated at constant pressure from 0oC to 100oC then work done will be

Grade:11

3 Answers

Vikas TU
7 years ago
W = nR(T2 – T1) => 1*R*(100 – 0) => 100R J would be the work done for One mole of an ideal gas when heated at constant pressure.
Daniel Livingston
6 years ago
W=-P.∆v =-P(V2-V1) =-P{(nRT2÷P)-(nRT1÷P)} =-nR(T2-T1) =-1×1.987(373-273) =-1.987(100) =-198.7 cal =-198.7×4.184 Joules =-831.3608 Joules
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
W = – P.∆V
= – P(V2 – V1)
= – P{(nRT2÷P) – (nRT1÷P)}
= – nR(T2 – T1)
= – 1×1.987(373 – 273)
= – 1.987(100)
= – 198.7 cal
= – 198.7 × 4.184 Joules
= – 831.3608 Joules
 
Thanks and regards,
Kushagra